OpenStudy (anonymous):

The area of a rectangular pool is given by the trinomial 8y^2 + 4y – 12. What are the possible dimensions of the pool? Use factoring.

OpenStudy (anonymous):

we need to find the length and width of the pool. there are many possibilities. so first factor completely. 8y^2 + 4y - 12 =4 * (2y^2 + y - 3) =4 * (2y + 3) * (y - 1) from here you can distribute the 4 to the other two factors. for example: (8y+12) by (y-1) could be a length X width. also, (2y+3) by (4y-4) could also be another. you can also split up the 4 as 2 x 2. so how many L X W can you come up with? I came up with 5.

OpenStudy (anonymous):

well, to find the possible measure of the dimensions of the pool using factoring, we need to equate the equation to zero.